Show that norm is continuous function.
1
Expert's answer
2020-05-01T18:50:04-0400

EXPLANATION.

Denote by "\left( X,\left\| \right\| \right)" the normalized space. "f(x)=\left\| x \right\| , \ f:X\rightarrow { R }_{ 1 }"

By the properties of the norm for any "x,a\in X" the inequalities"\left\| x \right\| =||x-a+a||\le \left\| x-a \right\| +\left\| a \right\|" , "\left\| a \right\| =||a-x+x||\le \left\| x-a \right\| +\left\| x \right\|" are true. Hence, "-\left( \left\| x \right\| -\left\| a \right\| \right) \le \left\| x-a \right\| \le \left( \left\| x \right\| -\left\| a \right\| \right)" , or

"\left| \left\| x \right\| -\left\| a \right\| \right| \le \left\| x-a \right\| \quad \quad". For the function "f" the last inequality means

"\left| f(x)-f(a) \right| \le \left\| x-a \right\|" (1)

Let "\u03b5>0, \u03b4=\u03b5" . Then from inequality (1) we obtain ,that for all "x,a\in X"

"\left\| x-a \right\| <\delta" implies "\left| f(x)-f(a) \right| \le \left\| x-a \right\| <\delta =ε" (2)

By the definition of continuous function ,(2) means, that "f" is continuous at any point "a\in X" .